# Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study over Open-ended Question Answering

### v2

@@ Abstract
Recent works integrating Knowledge Graphs (KGs) have led to promising improvements in enhancing the reasoning accuracy of Large Language Models (LLMs). However, current benchmarks focus mainly on closed-ended tasks, leaving a gap in the assessment of more complex real-world scenarios. This gap has also obscured the evaluation of KGs’ potential to mitigate the problem of hallucination in LLMs. To fill the gap, we introduce OKGQA, a new benchmark specifically designed to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions from different types, and incorporates specific metrics to measure both hallucination ratio and the enhancement in reasoning capabilities. To consider the scenario in which KGs may have varying levels of mistakes, we propose another benchmark variant OKGQA-P to assess model performance when the semantics and structure of KGs are deliberately perturbed and contaminated. OKGQA aims to (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design. We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs to reduce hallucination. The code and data are released at link.
@@ 1 Introduction
Contemporary LLMs are prone to producing hallucinations due to gaps in their knowledge (Gekhman et al., 2024; Lee et al., 2023), primarily when the training data contain misinformation, biases, or inaccuracies. These flaws can arise when we simply maximize the log-likelihood, leading to responses that may seem plausible, but are often irrelevant or incorrect (Weng, 2024). This issue is especially concerning in scenarios where accuracy and reliability are essential, such as healthcare (He et al., 2023) and science (Taylor et al., 2022).
@@
To address this critical limitation, researchers have turned to leveraging external knowledge graphs (KGs) as a complementary (Pan et al., 2024; Luo et al., 2023a; Hu et al., 2023; Sun et al., 2023; van Breugel et al., 2021). KGs provide several key advantages: they contain structured and explicit factual knowledge, including domain-specific information, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023). Furthermore, KGs offer traceability, allowing each piece of information to be traced back to its source. This traceability enables both verification of generated context and transparency in reasoning paths, making KGs a promising source for improving the reliability of LLM reasoning. Find more details of related works in §B.
However, current benchmarks for testing the capabilities of these LLM+KG models are predominantly closed-ended, restricting responses to a limited set of entities or relations (Talmor et al., 2019; Mihaylov et al., 2018; Jin et al., 2020) or a set of logical forms (Yih et al., 2016; Talmor and Berant, 2018; Puerto et al., 2023) derived from specific facts of KG (as shown in Figure 1). These benchmarks, while useful for evaluating basic retrieval and reasoning, are insufficient for testing the LLMs’ ability for faithful generation (i.e., minimize hallucination in LLM responses), leaving a gap in the assessment of complex, real-world scenarios.
Why Open-ended Question? A key limitation of closed-ended benchmarks is: they cannot assess the phenomenon of “more talk, more mistakes”, where longer, more complex responses tend to increase hallucination Qiu et al. (2024). Unlike closed-ended benchmarks that require short, specific answers, open-ended questions demand more detailed and nuanced explanations. This format also enables direct evaluation metrics like FActScore (Min et al., 2023) and SAFE (Wei et al., 2024), which measure factual consistency of LLM responses by decomposing them into atomic statements for comparison with external information. However, for short, rigid responses, this decomposition is not feasible, rendering these metrics inapplicable. Moreover, open-ended evaluations align more closely with human cognition Lu et al. (2024); Amirizaniani et al. (2024), where context-rich and interpretative answers are expected compared to simplistic and mechanical responses. Find more details of intuition in Appendix §A.1.
Motivation: Recognizing the research gap, our study aims to contribute to the existing knowledge from two perspectives: 1) explore whether KGs can make LLMs more trustworthy in the open-ended setting, and 2) conduct a comparative analysis to shed light on methods and directions for leveraging KGs to reduce LLMs’ hallucination.
@@
To achieve this, we develop a new benchmark, Open-ended Knowledge-Graphs Question Answering (OKGQA), specifically designed to assess LLMs enhanced with KGs under open-ended question answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions of different types and incorporates specific metrics to measure both the reduction in hallucinations and enhancements in reasoning capabilities. Specifically, all queries in OKGQA are synthesized by LLMs with pre-defined templates mentioned in Table 1 and corresponding Wikipedia context. To consider the scenarios in which KGs may have varying levels of mistakes (i.e., attributes may be mislabeled, relations may not exist, etc.), we propose another experiment setup OKGQA-P (§2.2) to assess model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated. We assess the hallucination ratio and the quality of the response (detailed metrics can be found in §4.1) of the tested models.
@@
Based on our experiments, we find that (1) retrieving KG information can mitigate factual errors in LLMs’ responses, especially for queries requiring additional reasoning; (2) directly performing reasoning in the LLM itself (e.g., internal reasoning strategies like Chain-of-thoughts (Kim et al., 2023) and self-consistency (Wang et al., 2022)) may cause bias and hallucination; (3) Subgraphs generally achieve the best performance across different query types, especially for simpler types of queries; (4) Integrating KGs can effectively reduce hallucinations in LLMs, even when KGs are contaminated to a certain degree.
@@
2 OKGQA: A Open-ended Knowledge Graph Question-Answering Benchmark
Overview. OKGQA is a comprehensive benchmark designed to assess how effectively LLMs enhanced with KGs perform in open-ended, real-world-like question answering scenarios. Unlike existing benchmarks that focus primarily on closed-ended tasks, OKGQA presents diverse open-ended question types that mirror the variable nature of practical applications. The benchmark evaluates how effectively models leverage KGs to generate both reliable and comprehensive answers, focusing on reducing hallucinations while maintaining reasoning capabilities of LLMs. To assess robustness under imperfect conditions, we also introduce OKGQA-P, a variant where KGs are deliberately perturbed and contaminated.
@@ 2.1 Dataset Construction
Total number of queries 850
@@
Queries. We utilize a template-based methodology to generate a diverse range of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries. Details regarding specific templates and example queries can be found in Table 1, while the corresponding prompts are provided in the Appendix C. To ensure that the generated queries represent real-world scenarios and complexities, we employ an iterative optimization approach that utilizes both automated and human evaluation to refine the query generation process (the details are given in Appendix A.2). Initially, we generate a diverse set of query candidates from a seed instruction. These candidates undergo automated evaluation using an LLM-based evaluator, which assigns quality scores $s_{\text{auto}}$ on a scale of 1-10, with higher scores indicating better performance across multiple metrics. Subsequently, human evaluators assess the same queries, producing corresponding normalized scores $s_{\text{human}}$ within the same range. To refine the dataset, we iteratively optimize the input instructions by minimizing the discrepancy between $s_{\text{human}}$ and $s_{\text{auto}}$. This optimization process ensures alignment between automated and human quality assessments. The queries are also categorized by complexity with detailed statistics in Figure 2.
KG Sub-graphs. To reduce the size of KGs, following previous works (Yih et al., 2016; Talmor and Berant, 2018), we sample subgraphs from DBpedia Knowledge Graph (Noted that all the queries in OKGQA can be reasoned based on DBPedia) by extracting all triples contained within the k-hop neighbors from the question entities in the query. We set k = 2 to balance graph size and computational feasibility, as increasing k leads to exponential growth in the number of edges and nodes, which can introduce excessive noise and make it more challenging to retrieve essential information from KGs [^1] This choice is also informed by common practices in other benchmarks, such as WebQSP Yih et al. (2016) and CWQ Talmor and Berant (2018), where 2-hop subgraphs are widely used for similar KGQA tasks.. To further reduce the size of the 2-hop subgraphs, we leverage personalized page-rank (PPR) (Bahmani et al., 2010) to prune the nodes/edges that are not relevant to the query (the details of the PPR algorithm are discussed in Appendix A.3). We compare the statistics of subgraphs before and after PPR pruning in Figure 2(a).
@@ 2.2 OKGQA-P: Benchmark with Noise & Perturbations in KGs
To further mimic the real situations where KGs may not be of high quality (i.e., attributes of nodes/edges may be mislabeled, relations may not exist, etc.), we propose another experiment setting OKGQA-P to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated. We introduce various perturbation scenarios including mislabeled attributes, incorrect relations, and missing connections to test the model’s reasoning capabilities with flawed or incomplete KG data. To quantify the degree of perturbation, we also evaluate both the semantic and structural similarity between the original and the modified KG as defined below.
@@
For the perturbation methods, we consider four perturbation heuristics based on (Raman et al., 2020) as follows: Relation Swapping (RS) randomly chooses two edges from $\mathcal{T}$ and swaps their relations. Relation Replacement (RR) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, and replaces the $e_{1}$ with another relation $e_{2}=\text{argmin}_{e\in\mathcal{E}}S_{\mathcal{G}}(v_{1},e,v_{2})$, where $S_{\mathcal{G}}(\cdot)$ is a KG score function adapted from ATS. The replacement is performed such that the resulting $(v_{1},e_{2},v_{2})$ has the closest score to the original triplet $(v_{1},e_{1},v_{2})$ to generate “harder negatives” - triplets that are semantically similar but incorrect, thereby challenging the model to different subtle distinctions. Noted that $S_{G}$ differs from the original ATS, as it operates on individual triplets rather than comparing two entire KGs. Edge Rewiring (ER) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, then replaces $v_{2}$ with another entity $v_{3}\in\mathcal{E}\backslash\mathcal{N}_{1}(v_{1})$, where $\mathcal{N}_{1}(v_{1})$ represents the 1-hop neighborhood of $v_{1}$. Edge Deletion (ED) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$ and deletes it. We control the perturbation level based on the percentage of KG edges perturbed. See results in Figures 5 and 6.
@@ 3 Exploring KG-augmented framework for reducing hallucination
To explore whether KG-augmented approaches can mitigate LLMs’ hallucination, we investigate a range of methods and discover that they can be encapsulated within a unified framework as shown in Figure 3. Our framework follows the paradigm of retrieval augmented generation (RAG) (Edge et al., 2024; Baek et al., 2023a), which retrieves essential information from the KGs, and then uses the retrieved knowledge to enhance the LLM’s generation. It has two components, i.e., Graph-guided retrieval (G-retrieval) and Graph-guided generation (G-generator), with a variety of algorithmic design choices. We analyze our proposed strategies within each component in §4, aiming to shed light on the best practices for leveraging KGs for reducing hallucinations in LLMs.
@@
Here we formalize the problem of context-relevant KG retrieval for open-ended question answering. Given a user query $q$, a pretrained language model (PLM) first encodes the input tokens, and models a conditional distribution $p(a|q)$ to generate an output response $a=[a_{1},\ldots,a_{T}]$. To explore whether KGs help reduce hallucinations of LLMs, we introduce the retrieved knowledge from the KG, denoted $\mathcal{Z}\in\mathcal{G}$, which can take various forms (e.g., triplets, paths, subgraphs, as shown in Figure 3). This retrieved knowledge contains information relevant to the user query, and either benefits the LLMs at its knowledge boundaries (Baek et al., 2023b) or explicitly guides its reasoning process (Luo et al., 2023a; Sun et al., 2023). The distribution of the retrieved knowledge $\mathcal{Z}$ is modeled as $p(\mathcal{Z}|q,\mathcal{G})$, leading to a new likelihood for generating responses, expressed as $p(a|q)$.
$p(a|q)$ $=\sum_{\mathcal{Z}\subseteq\mathcal{G}}p_{\phi}(a|q,\mathcal{Z})p_{\theta}(\mathcal{Z}|q,\mathcal{G})$ (1)
$\approx p_{\phi}(a|q,\mathcal{Z}^{*})p_{\theta}(\mathcal{Z}^{*}|q,\mathcal{G}),$
where $\mathcal{Z}^{*}=\text{argmax}_{\mathcal{Z}\in\mathcal{G}}p_{\theta}(\mathcal{Z}|q,\mathcal{G})$ represents the optimal subset of retrieved knowledge from the KGs, and $p_{\phi}(a|q,\mathcal{Z}^{*})$ represents the likelihood generating the response given the query $q$ and the retrieved knowledge $\mathcal{Z}^{*}$, parameterized by $\phi$ and $\theta$. Due to the potential exponential growth in the number of possible retrieved knowledge with graph size, efficient approximation methods are required. Consequently, the first line of Eq. 1 is approximated using the most probable retrieved knowledge $\mathcal{Z}^{*}$.

### v3

@@ Abstract
Recent works integrating Knowledge Graphs (KGs) have led to promising improvements in enhancing the reasoning accuracy of Large Language Models (LLMs). However, current benchmarks focus mainly on closed-ended tasks, leaving a gap in the assessment of more complex real-world scenarios. This gap has also obscured the evaluation of KGs’ potential to mitigate the problem of hallucination in LLMs. To fill the gap, we introduce OKGQA, a new benchmark specifically designed to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions from different types, and incorporates specific metrics to measure both hallucination ratio and the enhancement in reasoning capabilities. To consider the scenario in which KGs may have varying levels of mistakes, we propose another benchmark variant OKGQA-P to assess model performance when the semantics and structure of KGs are deliberately perturbed and contaminated. OKGQA aims to (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design. We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs to reduce hallucination.
@@ 1 Introduction
Contemporary LLMs are prone to producing hallucinations due to gaps in their knowledge Gekhman et al. (2024); Lee et al. (2023). These inaccuracies commonly stem from misinformation, biases, or errors in the training data, and lead to responses that seem plausible but may be irrelevant or incorrect Weng (2024). This issue is particularly concerning in high-stakes contexts such as healthcare He et al. (2023) and science Taylor et al. (2022) [^1] Code and data are released at https://anonymous.4open.science/r/OKGQA-CBB0.
@@
To address this limitation, researchers have turned to leveraging external knowledge graphs (KGs) as a complementary Pan et al. (2024); Luo et al. (2023a); Hu et al. (2023); Sui et al. (2024). KGs offer structured and explicit factual information—often domain-specific—and allow each piece of data to be traced back to its source Zheng et al. (2023); Agrawal et al. (2023). This traceability not only enables verification of the model’s reasoning but also brings transparency to the decision-making process, making KGs a promising method for enhancing the reliability of LLM reasoning. Find more details of related works in §C.
However, current benchmarks for testing these LLM+KG models are predominantly closed-ended Jin et al. (2020); Puerto et al. (2023), restricting the model’s output to a limited set of entities, relations Talmor et al. (2019); Mihaylov et al. (2018) or logical forms Yih et al. (2016); Talmor and Berant (2018). While these benchmarks are useful to measure retrieval and basic reasoning, they do not adequately capture whether a model is hallucinating. In closed-ended settings, errors can stem from incorrect retrieval or from fabricating (hallucinating) answers, yet conventional metrics (e.g., accuracy or precision) cannot distinguish between these two issues. This becomes problematic for more complex, real-world applications that demand nuanced answers Kantharaj et al. (2022).
In contrast, our work focuses on open-ended KGQA, where LLMs are prompted to generate more elaborate answers, include reasoning paths and supporting facts derived from the KG (as shown in Figure 1). This broader output space offers two key advantages: First, it enables direct assessments of hallucination with metrics like FActScore Min et al. (2023) or SAFE Wei et al. (2024)), which decompose longer responses into atomic statements for factual consistency checks with external knowledge sources like Wikipedia. Second, it increases the likelihood of exposing factual errors which helps to assess the phenomenon indicated in Qiu et al. (2024): where longer, more complex responses provide more opportunities for errors to occur. By adopting this open-ended approach, we aim to (1) explore whether KGs can make LLMs more trustworthy in the open-ended setting, and (2) conduct a comparative analysis to shed light on methods design and direction for leveraging KGs to reduce LLMs’ hallucination.
@@
To achieve this, we introduce a new benchmark, Open-ended Knowledge-Graphs Question Answering (OKGQA), specifically designed to assess LLMs enhanced with KGs in an open-ended QA setting. OKGQA is designed to closely reflect the complexities of practical applications with diverse questions types as mentioned in Table 1, ensuring that all the queries cannot be answered simply by retrieving isolated KG facts. To consider the scenarios for potentially contaminated or imperfect KGs (i.e., attributes may be mislabeled, relations may not exist, etc.), we also propose a variant OKGQA-P (§2.2) to assess model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated. In both settings, we measure the degree of hallucination and the overall quality of the responses (see §4.1 for details).
@@
Based on our experiments, we find that (1) integrating KG information generally mitigates factual errors, especially for queries requiring deep reasoning; (2) directly performing reasoning in the LLM itself (e.g., internal reasoning strategies like Chain-of-thought Kim et al. (2023) and Self-Consistency Wang et al. (2022)) may cause bias and hallucination; (3) subgraph-based methods often achieve the best performance for simpler query types; and (4) incorporating KGs effectively reduces hallucinations in LLMs even when the KG is partially contaminated.
@@
2 OKGQA: An Open-ended Knowledge Graph Question-Answering Benchmark
OKGQA is a comprehensive benchmark designed to assess how effectively LLMs enhanced with KGs perform in open-ended, real-world-like question answering scenarios. Unlike existing benchmarks that focus primarily on closed-ended tasks, OKGQA presents diverse open-ended question types that mirror the variable nature of practical applications. As illustrated in Figure 1, given a complex query and its corresponding subgraph in a KG, the system must be capable of understanding the relationships within the data and performing human-like reasoning over the KG content to compose a paragraph-long answer. In the following section, we first describe our dataset construction, including query generation via LLM templates and KG subgraph extraction with PPR pruning. We then introduce OKGQA-P, a benchmark variant designed to evaluate model robustness under KG perturbations, detailing our perturbation methods and the metrics used to assess semantic and structural deviations. Due to the page limitation, we also include some extension of our benchmark in Appendix §B, including multilingual setup, and more analysis.
@@ 2.1 Dataset Construction
Total number of queries 850 $\rightarrow$ 2,050
@@
Queries.
We utilize a template-based approach to generate a diverse range of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries. Details regarding specific templates and example queries can be found in Table 1, while the corresponding prompts are provided in the Appendix D. To ensure that the generated queries represent real-world scenarios and complexities, we employ an iterative optimization approach that utilizes both automated and human evaluation to refine the query generation process (the details are given in Appendix A.1). Initially, we generate a diverse set of query candidates from a seed instruction. These candidates undergo automated evaluation using an LLM-based evaluator, which assigns quality scores $s_{\text{auto}}$ on a scale of 1-10, with higher scores indicating better performance across multiple metrics. Subsequently, human evaluators assess the same queries, producing corresponding normalized scores $s_{\text{human}}$ within the same range. To refine the dataset, we iteratively optimize the input instructions by minimizing the discrepancy between $s_{\text{human}}$ and $s_{\text{auto}}$. This optimization process ensures alignment between automated and human quality assessments. The queries are also categorized by complexity with detailed statistics in Figure 2.
KG Sub-graphs.
To reduce the size of KGs while retaining relevant information, we follow previous work Yih et al. (2016); Talmor and Berant (2018) by sampling subgraphs from DBpedia (Noted that all queries in OKGQA can be answered using DBpedia). We extract all triples contained within the $K$-hop neighbors from the entities mentioned in the query. We set $K=2$ to balance graph size and computational feasibility. As increasing beyond $2$-hop subgraphs generally leads to exponential growth in edges and nodes Jin et al. (2020), which increase excessive noise and complicating information retrieval [^2] This choice is also informed by common practices in other benchmarks, such as WebQSP Yih et al. (2016) and CWQ Talmor and Berant (2018), where 2-hop subgraphs are widely used for similar KGQA tasks.. To further reduce the size of the 2-hop subgraphs, we leverage Personalized Page-Rank (PPR) Bahmani et al. (2010) to prune the nodes/edges that are not relevant to the query (the details of the PPR algorithm are discussed in Appendix A.2). We compare the statistics of subgraphs before and after PPR pruning in Figure 2(a).
@@ 2.2 OKGQA-P: Benchmark with Noise & Perturbations in KGs
KGs are often annotated by humans and can contain errors such as mislabeled attributes or missing relations. To mimic the real situations where KGs’ quality may not be fully reliable, we propose OKGQA-P to assess the model performance under deliberately perturbed and contaminated KGs. We introduce various perturbation scenarios including mislabeled attributes, incorrect relations, and missing connections to test how well models can handle flawed or incomplete KG data. To quantify the degree of perturbation, we measure the semantic and structural similarity between the original and the modified KG as defined below.
@@
For the perturbation methods, we consider four edge-based perturbation heuristics based on Raman et al. (2020) as follows:
  • Relation Swapping (RS) randomly chooses two edges from $\mathcal{T}$ and swaps their relations.
  • Relation Replacement (RR) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, and replaces the $e_{1}$ with another relation $e_{2}=\mathrm{argmin}_{e\in\mathcal{E}}S_{\mathcal{G}}(v_{1},e,v_{2})$, where $S_{\mathcal{G}}(\cdot)$ is a KG score function adapted from ATS. This yield “harder negatives” - triplets that are semantically similar but incorrect.
  • Edge Rewiring (ER) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, then replaces $v_{2}$ with another entity $v_{3}\in\mathcal{E}\backslash\mathcal{N}_{1}(v_{1})$, where $\mathcal{N}_{1}(v_{1})$ represents the 1-hop neighborhood of $v_{1}$.
  • Edge Deletion (ED) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$ and deletes it.
We control perturbation level by adjusting the percentage of edges in $\mathcal{G}$ that are perturbed. Refer to Figures 5 and 6 for empirical results.
@@ 3 Exploring KG-augmented framework for Reducing Hallucination
To explore whether KG-augmented approaches can mitigate LLMs’ hallucination, we propose a unified framework as shown in Figure 3. Our framework follows the paradigm of retrieval augmented generation (RAG) Edge et al. (2024); Baek et al. (2023a), which retrieves essential information from the KGs, and then uses the retrieved knowledge to enhance the LLM’s generation (§3.1). It consists of two components, i.e., Graph-guided retrieval3.2) and Graph-guided generator3.3), with a variety of algorithmic design choices. We analyze the strategies within each component in §4, aiming to shed light on the best practices for leveraging KGs for reducing hallucinations in LLMs.
3.1 Formalization
We formalize the KG-augmented framework for reducing hallucination as follows. Given a user query $q$, a pretrained language model generates a paragraph-like answer $a$ by modeling the conditional probability $p(a|q)$. To explore whether KGs help reduce hallucinations of LLMs, we introduce retrieved knowledge $\mathcal{Z}$ from the KG and define:
$p(a|q)=\sum_{\mathcal{Z}\subseteq\mathcal{G}}p_{\phi}(a|q,\mathcal{Z})p_{\theta}(\mathcal{Z}|q,\mathcal{G}),$ (1)
where $p_{\phi}(a|q,\mathcal{Z})$ is the likelihood of generating the paragraph-like answer $a$ conditioned on $q$ and $\mathcal{Z}$ (parameterized by $\phi$), and $p_{\theta}(\mathcal{Z}|q,\mathcal{G})$ models the retrieval of knowledge subsets (parameterized by $\theta$). Because the number of possible subsets $\mathcal{Z}$ can be exponentially large relative to the size of $\mathcal{G}$, we approximate the sum by selecting the most probable knowledge subset: $\mathcal{Z}^{*}=\text{argmax}_{\mathcal{Z}\in\mathcal{G}}p_{\theta}(\mathcal{Z}|q,\mathcal{G})$, yielding:
$p(a|q)\approx p_{\phi}(a|q,\mathcal{Z}^{*})p_{\theta}(\mathcal{Z}^{*}|q,\mathcal{G})$ (2)